OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? a+98=-2

jimthompson5910 (jim_thompson5910):

What can you add to +98 to get 0?

OpenStudy (anonymous):

-98

jimthompson5910 (jim_thompson5910):

Good, so do this to both sides a+98=-2 a+98 + (-98) = -2 + (-98) a + 0 = -100 a = -100